%% BioMed_Central_Bib_Style_v1.01

\begin{thebibliography}{19}
% BibTex style file: bmc-mathphys.bst (version 2.1), 2014-07-24
\ifx \bisbn   \undefined \def \bisbn  #1{ISBN #1}\fi
\ifx \binits  \undefined \def \binits#1{#1}\fi
\ifx \bauthor  \undefined \def \bauthor#1{#1}\fi
\ifx \batitle  \undefined \def \batitle#1{#1}\fi
\ifx \bjtitle  \undefined \def \bjtitle#1{#1}\fi
\ifx \bvolume  \undefined \def \bvolume#1{\textbf{#1}}\fi
\ifx \byear  \undefined \def \byear#1{#1}\fi
\ifx \bissue  \undefined \def \bissue#1{#1}\fi
\ifx \bfpage  \undefined \def \bfpage#1{#1}\fi
\ifx \blpage  \undefined \def \blpage #1{#1}\fi
\ifx \burl  \undefined \def \burl#1{\textsf{#1}}\fi
\ifx \doiurl  \undefined \def \doiurl#1{\url{https://doi.org/#1}}\fi
\ifx \betal  \undefined \def \betal{\textit{et al.}}\fi
\ifx \binstitute  \undefined \def \binstitute#1{#1}\fi
\ifx \binstitutionaled  \undefined \def \binstitutionaled#1{#1}\fi
\ifx \bctitle  \undefined \def \bctitle#1{#1}\fi
\ifx \beditor  \undefined \def \beditor#1{#1}\fi
\ifx \bpublisher  \undefined \def \bpublisher#1{#1}\fi
\ifx \bbtitle  \undefined \def \bbtitle#1{#1}\fi
\ifx \bedition  \undefined \def \bedition#1{#1}\fi
\ifx \bseriesno  \undefined \def \bseriesno#1{#1}\fi
\ifx \blocation  \undefined \def \blocation#1{#1}\fi
\ifx \bsertitle  \undefined \def \bsertitle#1{#1}\fi
\ifx \bsnm \undefined \def \bsnm#1{#1}\fi
\ifx \bsuffix \undefined \def \bsuffix#1{#1}\fi
\ifx \bparticle \undefined \def \bparticle#1{#1}\fi
\ifx \barticle \undefined \def \barticle#1{#1}\fi
\bibcommenthead
\ifx \bconfdate \undefined \def \bconfdate #1{#1}\fi
\ifx \botherref \undefined \def \botherref #1{#1}\fi
\ifx \url \undefined \def \url#1{\textsf{#1}}\fi
\ifx \bchapter \undefined \def \bchapter#1{#1}\fi
\ifx \bbook \undefined \def \bbook#1{#1}\fi
\ifx \bcomment \undefined \def \bcomment#1{#1}\fi
\ifx \oauthor \undefined \def \oauthor#1{#1}\fi
\ifx \citeauthoryear \undefined \def \citeauthoryear#1{#1}\fi
\ifx \endbibitem  \undefined \def \endbibitem {}\fi
\ifx \bconflocation  \undefined \def \bconflocation#1{#1}\fi
\ifx \arxivurl  \undefined \def \arxivurl#1{\textsf{#1}}\fi
\csname PreBibitemsHook\endcsname

%%% 1
\bibitem[\protect\citeauthoryear{Noys}{2010}]{noys2010persistence}
\begin{botherref}
\oauthor{\bsnm{Noys}, \binits{B.}}:
Persistence of the negative: A critique of contemporary continental theory.
Edinburgh University Press
(2010)
\end{botherref}
\endbibitem

%%% 2
\bibitem[\protect\citeauthoryear{Licata et~al.}{2017}]{licata2017fibrational}
\begin{bchapter}
\bauthor{\bsnm{Licata}, \binits{D.R.}},
\bauthor{\bsnm{Shulman}, \binits{M.}},
\bauthor{\bsnm{Riley}, \binits{M.}}:
\bctitle{A fibrational framework for substructural and modal logics}.
In: \bbtitle{2nd International Conference on Formal Structures for Computation
  and Deduction (FSCD 2017)}
(\byear{2017}).
\bcomment{Schloss Dagstuhl-Leibniz-Zentrum fuer Informatik}
\end{bchapter}
\endbibitem

%%% 3
\bibitem[\protect\citeauthoryear{de~Paiva and Ritter}{2016}]{de2016fibrational}
\begin{barticle}
\bauthor{\bsnm{Paiva}, \binits{V.}},
\bauthor{\bsnm{Ritter}, \binits{E.}}:
\batitle{Fibrational modal type theory}.
\bjtitle{Electronic Notes in Theoretical Computer Science}
\bvolume{323},
\bfpage{143}--\blpage{161}
(\byear{2016})
\end{barticle}
\endbibitem

%%% 4
\bibitem[\protect\citeauthoryear{Gratzer et~al.}{2020}]{gratzer2020multimodal}
\begin{bchapter}
\bauthor{\bsnm{Gratzer}, \binits{D.}},
\bauthor{\bsnm{Kavvos}, \binits{G.}},
\bauthor{\bsnm{Nuyts}, \binits{A.}},
\bauthor{\bsnm{Birkedal}, \binits{L.}}:
\bctitle{Multimodal dependent type theory}.
In: \bbtitle{Proceedings of the 35th Annual ACM/IEEE Symposium on Logic in
  Computer Science},
pp. \bfpage{492}--\blpage{506}
(\byear{2020})
\end{bchapter}
\endbibitem

%%% 5
\bibitem[\protect\citeauthoryear{Awodey}{2018}]{awodey2018univalence}
\begin{barticle}
\bauthor{\bsnm{Awodey}, \binits{S.}}:
\batitle{Univalence as a principle of logic}.
\bjtitle{Indagationes Mathematicae}
\bvolume{29}(\bissue{6}),
\bfpage{1497}--\blpage{1510}
(\byear{2018})
\end{barticle}
\endbibitem

%%% 6
\bibitem[\protect\citeauthoryear{Pastro}{2010}]{pastro2010note}
\begin{botherref}
\oauthor{\bsnm{Pastro}, \binits{C.}}:
Note on star-autonomous comonads.
arXiv preprint arXiv:1010.5304
(2010)
\end{botherref}
\endbibitem

%%% 7
\bibitem[\protect\citeauthoryear{Cockett et~al.}{2006}]{cockett2006coherence}
\begin{barticle}
\bauthor{\bsnm{Cockett}, \binits{J.}},
\bauthor{\bsnm{Hasegawa}, \binits{M.}},
\bauthor{\bsnm{Seely}, \binits{R.}}:
\batitle{Coherence of the double involution on $\star-autonomous$ categories}.
\bjtitle{Theory and Applications of Categories}
\bvolume{17}(\bissue{2}),
\bfpage{17}--\blpage{29}
(\byear{2006})
\end{barticle}
\endbibitem

%%% 8
\bibitem[\protect\citeauthoryear{Boyarchenko and
  Drinfeld}{2013}]{boyarchenko2013duality}
\begin{barticle}
\bauthor{\bsnm{Boyarchenko}, \binits{M.}},
\bauthor{\bsnm{Drinfeld}, \binits{V.}}:
\batitle{A duality formalism in the spirit of grothendieck and verdier}.
\bjtitle{Quantum Topology}
\bvolume{4}(\bissue{4}),
\bfpage{447}--\blpage{489}
(\byear{2013})
\end{barticle}
\endbibitem

%%% 9
\bibitem[\protect\citeauthoryear{Abramsky et~al.}{2017}]{abramsky2017pebbling}
\begin{bchapter}
\bauthor{\bsnm{Abramsky}, \binits{S.}},
\bauthor{\bsnm{Dawar}, \binits{A.}},
\bauthor{\bsnm{Wang}, \binits{P.}}:
\bctitle{The pebbling comonad in finite model theory}.
In: \bbtitle{2017 32nd Annual ACM/IEEE Symposium on Logic in Computer Science
  (LICS)},
pp. \bfpage{1}--\blpage{12}
(\byear{2017}).
\bcomment{IEEE}
\end{bchapter}
\endbibitem

%%% 10
\bibitem[\protect\citeauthoryear{Abramsky and
  Marsden}{2021}]{abramsky2021comonadic}
\begin{bchapter}
\bauthor{\bsnm{Abramsky}, \binits{S.}},
\bauthor{\bsnm{Marsden}, \binits{D.}}:
\bctitle{Comonadic semantics for guarded fragments}.
In: \bbtitle{2021 36th Annual ACM/IEEE Symposium on Logic in Computer Science
  (LICS)},
pp. \bfpage{1}--\blpage{13}
(\byear{2021}).
\bcomment{IEEE}
\end{bchapter}
\endbibitem

%%% 11
\bibitem[\protect\citeauthoryear{Abramsky and
  Marsden}{2022}]{abramsky2022comonadic}
\begin{bchapter}
\bauthor{\bsnm{Abramsky}, \binits{S.}},
\bauthor{\bsnm{Marsden}, \binits{D.}}:
\bctitle{Comonadic semantics for hybrid logic}.
In: \bbtitle{47th International Symposium on Mathematical Foundations of
  Computer Science (MFCS 2022)}
(\byear{2022}).
\bcomment{Schloss Dagstuhl-Leibniz-Zentrum f{\"u}r Informatik}
\end{bchapter}
\endbibitem

%%% 12
\bibitem[\protect\citeauthoryear{Postniece}{2010}]{postniece2010proof}
\begin{botherref}
\oauthor{\bsnm{Postniece}, \binits{L.}}:
Proof theory and proof search of bi-intuitionistic and tense logic.
PhD thesis,
Australian National University
(2010)
\end{botherref}
\endbibitem

%%% 13
\bibitem[\protect\citeauthoryear{Onishi}{2015}]{onishi2015substructural}
\begin{botherref}
\oauthor{\bsnm{Onishi}, \binits{T.}}:
Substructural negations.
The Australasian Journal of Logic
\textbf{12}(4)
(2015)
\end{botherref}
\endbibitem

%%% 14
\bibitem[\protect\citeauthoryear{Lahav et~al.}{2017}]{lahav2017sequent}
\begin{barticle}
\bauthor{\bsnm{Lahav}, \binits{O.}},
\bauthor{\bsnm{Marcos}, \binits{J.}},
\bauthor{\bsnm{Zohar}, \binits{Y.}}:
\batitle{Sequent systems for negative modalities}.
\bjtitle{Logica Universalis}
\bvolume{11},
\bfpage{345}--\blpage{382}
(\byear{2017})
\end{barticle}
\endbibitem

%%% 15
\bibitem[\protect\citeauthoryear{Ciabattoni
  et~al.}{2014}]{ciabattoni2014hypersequent}
\begin{barticle}
\bauthor{\bsnm{Ciabattoni}, \binits{A.}},
\bauthor{\bsnm{Ramanayake}, \binits{R.}},
\bauthor{\bsnm{Wansing}, \binits{H.}}:
\batitle{Hypersequent and display calculi--a unified perspective}.
\bjtitle{Studia Logica}
\bvolume{102},
\bfpage{1245}--\blpage{1294}
(\byear{2014})
\end{barticle}
\endbibitem

%%% 16
\bibitem[\protect\citeauthoryear{Odintsov and
  Wansing}{2021}]{odintsov2021routley}
\begin{barticle}
\bauthor{\bsnm{Odintsov}, \binits{S.}},
\bauthor{\bsnm{Wansing}, \binits{H.}}:
\batitle{Routley star and hyperintensionality}.
\bjtitle{Journal of Philosophical Logic}
\bvolume{50},
\bfpage{33}--\blpage{56}
(\byear{2021})
\end{barticle}
\endbibitem

%%% 17
\bibitem[\protect\citeauthoryear{Ahrens and
  Lumsdaine}{2019}]{ahrens2019displayed}
\begin{botherref}
\oauthor{\bsnm{Ahrens}, \binits{B.}},
\oauthor{\bsnm{Lumsdaine}, \binits{P.L.}}:
Displayed categories.
Logical Methods in Computer Science
\textbf{15}
(2019)
\end{botherref}
\endbibitem

%%% 18
\bibitem[\protect\citeauthoryear{Gehrke}{2014}]{gehrke2014canonical}
\begin{botherref}
\oauthor{\bsnm{Gehrke}, \binits{M.}}:
Canonical extensions, esakia spaces, and universal models.
Leo Esakia on duality in modal and intuitionistic logics,
9--41
(2014)
\end{botherref}
\endbibitem

%%% 19
\bibitem[\protect\citeauthoryear{Celani and Jansana}{2014}]{celani2014easkia}
\begin{botherref}
\oauthor{\bsnm{Celani}, \binits{S.A.}},
\oauthor{\bsnm{Jansana}, \binits{R.}}:
Easkia duality and its extensions.
Leo Esakia on Duality in Modal and Intuitionistic Logics,
63--98
(2014)
\end{botherref}
\endbibitem

\end{thebibliography}

\documentclass[sn-mathphys,Numbered,pdflatex]{sn-jnl}% Math and Physical Sciences Reference Style
\usepackage{graphicx}%
\usepackage{multirow}%
\usepackage{amsmath,amssymb,amsfonts}%
\usepackage{amsthm}%
\usepackage{mathrsfs}%
\usepackage[title]{appendix}%
\usepackage{xcolor}%
\usepackage{textcomp}%
\usepackage{manyfoot}%
\usepackage{lmodern}%
\usepackage{oplotsymbl}
\usepackage{booktabs}%
\usepackage{algorithm}%
\usepackage{algorithmicx}%
\usepackage{algpseudocode}%
\usepackage{listings}%
\newtheorem{theorem}{Theorem}%  meant for continuous numbers
\newtheorem{proposition}[theorem]{Proposition}% 

\begin{document}

\title[Article Title]{\centering Negation and Identity in a Modal Mode Theory \\or\\the Persistence of the Negative\footnote{An invocation of \cite{noys2010persistence}}}

%%=============================================================%%
%% Prefix	-> \pfx{Dr}
%% GivenName	-> \fnm{Joergen W.}
%% Particle	-> \spfx{van der} -> surname prefix
%% FamilyName	-> \sur{Ploeg}
%% Suffix	-> \sfx{IV}
%% NatureName	-> \tanm{Poet Laureate} -> Title after name
%% Degrees	-> \dgr{MSc, PhD}
%% \author*[1,2]{\pfx{Dr} \fnm{Joergen W.} \spfx{van der} \sur{Ploeg} \sfx{IV} \tanm{Poet Laureate} 
%%                 \dgr{MSc, PhD}}\email{iauthor@gmail.com}
%%=============================================================%%

\author{\fnm{Juan} \sur{Afanador}}\email{jafanad@gmail.com}

% \affil*[1]{\orgdiv{Department}, \orgname{Organization}, \orgaddress{\street{Street}, \city{City}, \postcode{100190}, \state{State}, \country{Country}}}

\keywords{Non-trivial Negation $\cdot$ Adjunctor Equivalence $\cdot$ Intuitionistic Modal Logic $\cdot$ Mode Theory}

%%\pacs[JEL Classification]{D8, H51}

%%\pacs[MSC Classification]{35A01, 65L10, 65L12, 65L20, 65L70}

% \date{31 August 2023}

\maketitle

\section*{}\label{sec1}
\vspace{-1cm}
This piece threads substructurality and modality into a negation that activates the downside of equivalence and identity in a fibrational framework \cite{licata2017fibrational, de2016fibrational, gratzer2020multimodal}. The piece is a working through of negation and contradiction as type-theoretic/categorial objects, towards an immanent critique of the subtending univalent paradigm (v. \cite{awodey2018univalence}). Although this is not the terminus of the piece, i wish to try and delineate the epistemic and intra-mundane problematics intertwined therewith. 
% -- sociation warrants that thinking subverts itself. 

The piece's terminus is a mode theory of an intuitionistic modal logic $\mathcal{L}_{nod}$ that internalises a restriction on the Double Negation Elimination (DNE) rule. The meta-negations of $\mathcal{L}_{nod}$ give way to a strictly weak $\star-autonomous$ \cite{pastro2010note, cockett2006coherence} (sc. Grothendieck-Verdier \cite{boyarchenko2013duality}) category, which warrants a comonadic characterisation of contradiction. The proposed comonadic semantics may be applied to model theory \cite{abramsky2017pebbling}, deductive complexity \cite{abramsky2021comonadic}, or combinatorics \cite{abramsky2022comonadic}, but it is in displaying  the tensions between negation (sc. the negative) and identity (sc. the affirmative), inherent in such instrumentalisms, where the relevance of the semantics may lie. 

To work my case, i use an intuitionistic modal language akin to $LBiKt$ under standard notation and Kripke semantics \cite{postniece2010proof}. That is, $\mathcal{M}=\langle W, \smallfrown, \smallsmile, \mathcal{I} \rangle$ is a model defined by the frame formed by a set of worlds $W$, world-accessibility relations $\smallfrown, \smallsmile\subseteq W\times W$ \cite{onishi2015substructural, lahav2017sequent}, and an interpretative mapping $\mathcal{I}:Atom(\mathcal{L}_{nod}) \mapsto 2^W$, whereupon satisfiability and validity are as habitually posed. $\mathcal{L}_{nod}$ extends $\mathcal{L}_{BiN}$ (v. \cite{ciabattoni2014hypersequent} and \cite{odintsov2021routley}) to accommodate the operators $\{\blacktriangleright, \blacktriangleleft, \triangleright, \triangleleft\}$, allowing for a modal (tense) split of negation -- e.g., $\mathcal{M}, w\vDash\blacktriangleright A,\, \exists v(v\smallsmile w\, \text{and}\, \mathcal{M},v\nvDash A)$ and $\mathcal{M}, w\vDash\triangleright A,\, \forall v(v\smallfrown w\, \text{then}\, \mathcal{M},v\nvDash A)$ for $w, v\in W$. (In)exhaustiveness as (the) failure (of the concept to model its object) is key.

Contexts in $\mathcal{L}_{nod}$ are of two types, either $X::=I|\alpha\in \mathcal{L}_{nod}| \sharp X|\flat X|X\pentagodot Y|X>Y$ or $\Gamma[\,]::=[\,]|\Gamma[\,]\pentagodot (X)| (X)\pentagodot \Gamma[\,]$. Following the $LBiInt_1$ usage, $X > \Gamma[\,]$ is a negative context. Structural and object-language operators can be mapped, with little overload, like so: $I\mapsto \{\top, \bot\}$, $\pentagodot\,\mapsto \{\land, \lor\}$, $>\,\mapsto \{-\hspace{-.8em}<, \to\}$, $\sharp\mapsto\{\blacktriangleright, \triangleright\}$, and $\flat\mapsto\{\blacktriangleleft, \triangleleft\}$. Exemplar rules of the attendant sequent calculus are $x:\flat A \vdash Y/x:\blacklozenge A \vdash Y$ and $\Gamma, F_{x\pentagodot y}(x:\blacklozenge A, y:\flat A)/\Gamma, x:\blacklozenge A$.

In our mode theory, frames index modes, $\{\sharp, \flat\}$ connote their modalities, and maps between them prefigure type-theoretic 2-cells. Alternatively, these are the objects, morphisms, and (categorical) 2-cells of the $\mathbf{Struct}$ 2-category \cite{abramsky2022comonadic}. In the forthcoming piece i lift the category these modes inhabit to a monadic (display) category (viz. \cite{ahrens2019displayed}), internalising the notion of contexts noted above to show that DNE must needs not be. Or less informally:

\begin{proposition} \label{prop}
There exists, at least, one mode $A\in \mathbf{Struct}$, such that the canonical morphism of the symmetric closed monoidal category $\langle \mathbf{Struct}, \pentagodot, I, \to \rangle$ is not an isomorphism.
\end{proposition}

To prove Proposition \ref{prop}, the dualising unit of the monoidal category is derived from the modal split of negation and the contextual interpretation of modalities. Although algebraically germane, the dualising techniques of Priestley \cite{gehrke2014canonical} and Esakia \cite{celani2014easkia} inform our derivation of duality. Once this result is established, the covariant functor described by the dualising unit serves to probe for modes which cannot be identified with their double negations. Or more eloquently, we are interested in the fragments of the nameless mode theory that Proposition \ref{prop} inhabits, where what is negated stays negated.

In closing i rehearse the opening caveat -- the subversion of (adjunctor) equivalence, in our mode-theoretic rendition of negation, also constitutes a point of entry into a questioning of the Univalence Axiom as a principle of logic \cite{awodey2018univalence}. In unfolding this piece two themes are obliquely interweaved: the metonymic claim of essential sameness that the univalent paradigm articulates in the computarised idiom of a post-theoretical world; and the adjacent abolition of the transcendental subject from logic (sc. cognition), whereby the intervening agency between individual action and social norm is subl(im)ated. Thinking negativity (sc. negation) may rehabilitate a critical engagement with the (social) aporias of what is. 

\bibliography{neging}% common bib file
%% if required, the content of .bbl file can be included here once bbl is generated
%%\input sn-article.bbl

\end{document}